# Advancing Superconducting Qubits: CMOS-Compatible Processing and Room Temperature Characterization for Scalable Quantum Computing beyond 2D Architectures

### v6

@@ Abstract
We report on an industry-grade CMOS-compatible qubit fabrication approach using a CMOS pilot line, enabling a yield of functional devices reaching 92.8 %, with a resistance spread evaluated across the full wafer 200 mm diameter of 12.4 % and relaxation times () approaching 80 µs. Furthermore, we conducted a comprehensive analysis of wafer-scale room temperature (RT) characteristics collected from multiple wafers and fabrication runs, focusing on RT measurements and their correlation to low temperature qubit parameters. From defined test structures, a across-wafer junction area variation of 10.1 % and oxide barrier variation of 7.2 % was calculated. Additionally, we show a close-correlation between qubit junction resistance and frequency in accordance with the Ambegaokar-Baratoff relation with a critical temperature 0.71 K. This relation sets the stage for pre-cooldown qubit evaluation and sorting. In particular, such early-on device characterization and validation are crucial for increasing the fabrication yield and qubit frequency targeting, which currently represent major scaling challenges. Furthermore, it enables the fabrication of large multichip quantum systems in the future. Our findings highlight the great potential of CMOS-compatible industry-style fabrication of superconducting qubits for scalable quantum computing in a foundry pilot line cleanroom.
@@ I Introduction
Scalable, CMOS-compatible processing on 200 mm wafers is a critical step toward industrializing superconducting qubit-based quantum processing units [1]. It enables the precise fabrication and complex 3D integration needed for advanced quantum architectures [2, 3]. By leveraging mature semiconductor manufacturing methods, this approach ensures reproducibility, scalability, and the transition from laboratory prototypes to large-scale quantum systems. The fabrication of CMOS-compatible superconducting all-aluminum junctions using a double dry-etch technique with industry-standard manufacturing tools has been successfully demonstrated [4, 5, 6]. In addition, comprehensive wafer-scale characterization at RT has been reported [7, 8]. A recent study [4] integrated both approaches, showcasing the benefits of industry-compatible processes for transmon-style qubit fabrication, achieving a high yield of 98.5 % and relaxation times () of up to 167 µs. The study indicated that junction area and oxide barrier are essential parameters for the precise frequency targeting. As they vary across the wafer, they are worth monitoring for improved device characterisation. Additionally, like others [9, 10], they used critical dimension scanning electron microscopy (CD-SEM) to evaluate their structuring process at RT. Subsequently, they analyzed selected qubits after cooldown in the cryostat for benchmarking. This low temperature characterization, which is often very time-consuming and performed at die-level after leaving the production line, has become an established method for qubit characterization [11]. Currently, there is significant research focused on developing time-efficient, wafer-scale characterization methods for qubits during fabrication as part of the process control monitoring (PCM). This approach rely on RT techniques, as traditional cryogenic methods are impractical for large-scale analysis. Simple resistance measurements on test structures, like shorts and test junctions, can be used for monitor the line width variation as well as the junction oxide barrier variation across the 200 mm wafer. Additionally, a selection for functional qubit chips can be implemented to increase the yield for the subsequent process steps, as non-functional devices can be excluded early on. The primary aim of this wafer-scale characterization is to predict the cryogenic behavior of qubit devices, including the qubit frequency targeting, based on measurements conducted at RT.
@@
This present study demonstrates a comprehensive methodology for the fabrication and RT characterization of superconducting qubits on a 200 mm silicon wafer, emphasizing the importance of detailed wafer-level analysis for process monitoring. By integrating an array of test structures, including test junctions and shorts, our approach facilitates the precise evaluation of junction characteristics at RT and variation across the entire 200 mm wafer. For example, the analysis of shorts enabled an efficient evaluation of the metal structuring variability in terms of the junction area variation without using CD-SEM. Measuring the junctions resistance enabled us to select for functional devices, which had a high yield across the full 200 mm wafer. In combination with the cryogenic analysis of those functional qubit chips, we could demonstrate validity of the Ambegaokar-Baratoff relation for all our samples over a large resistance range. This enables us to select suitable qubit chips based on their resistance already at RT at wafer-level. Among others, this approach is especially important for further processing in 3D integrated architectures for increased yield. Our study highlights the effectiveness of the proposed characterization principles for CMOS-compatible superconducting qubits, which are crucial for achieving time-efficient process control and optimization for scaling of a QPU.
@@ II Methods
In total 375 identical dies with a size of 10 mm x 7 mm containing three different sub-dies were fabricated on a single wafer (Fig. 1a). The relevant sub-die, shown in Fig. 1b, is the qubit chip with a single reference resonator and four fixed-frequency qubits with equal Josephson junction (JJ) area, all addressed through a common RF feedline. Additionally, a process control monitoring (PCM) array was placed in the wafer’s dicing street. It consist of 16 test junctions of sizes varying from 0.1225 µ up to 1 µ and eight shorts (i.e. simple metal lines with landing pads for the prober needles) with widths from 0.35 µm to 1 µm. The fixed-frequency qubits were used to analyze the cryogenic performance of the fabrication process by measuring the qubits transition frequencies (), relaxation times (), and dephasing times (). With the test junctions, RT resistance and IV measurements were performed to access junction’s resistances and breakthrough voltage without harming the qubits JJ. Additionally, the shorts were used to obtain information about the JJ electrode width distribution across the 200 mm wafer, which is limited by the structuring process. Both test structures were used to maintain a time-efficient RT PCM during fabrication.
@@ III.1 Shorts Analysis
In general, for all our Al depositions for the layer thicknesses of 150 nm and 100 nm we measured of 1.6 %. Therefore, we assumed this value for the BE and TE deposition for the samples in this paper. For this, we obtain a value of = 5.1 %, which is smaller than the BE width distribution of = 8.7 %. Such increase in width variability for decreasing design width is frequently observed for metal structuring [10]. Additionally, a = 10.1 % for the area variation is obtained from the relation
The , observed for our junctions, with an area of 0.175 µm2, is notably larger than compared to values () reported sizes for JJs with areas (A 0.04 µm2) typically used in superconducting qubits [4, 10, 9]. When using optical lithography, the design widths approach the limits of the i-line stepper (365 nm) tool, which causes an increase in area variability. We note that our time-efficient determination of electrode width variability, which applies only RT measurements of test structures, is much faster than the time-consuming optical inspection with critical dimension scanning electron microscope (CD-SEM) that is commonly used.
Furthermore, additional shorts with different widths can be used to determine the offset between the design width and actual width . As the across-wafer resistance mean values of the respective shorts is inverse proportional to , we can define . This approach can be used for both metal layers separately. The offset of the structuring process to the design value can be derived from Fig. 3b at the crossing of the fit with the y-axis and is = - 99.3 nm for the BE and = - 45.4 nm for the TE. These offsets derive from approaching the feature size limit of the optical lithography. Thus, our PCM of shorts enable not only the determination of the area size variability across the wafer () in a time-efficient manner, but also provides a calculation of the offset from the design widths .
@@ III.2 Junction analysis
= 7.2 % is smaller than the area spread of = 10.1 %, indicating that the contribution of the spread in oxide thickness and barrier height is not dominant, although it remains larger than in related works [4]. Therefore, we further analyzed our junctions in terms of their tunnel barrier characteristics by means of extended current voltage measurements. An example is shown in Fig. 5a. In general, the current through the junction can be described with a simple power-law dependence [15]
@@ IV.1 Frequency spectroscopy
To address qubits, read-out resonators were used for qubit characterization. By determining the “punch-out” power, i.e. the readout power at which the induced current exceeds the critical current of the junction, one-tone spectroscopy (Fig. 6a) serves as a first check whether an active qubit is coupled to the resonator. The strength of this coupling is indicated by magnitude of the shift. Two-tone spectroscopy (Fig. 6b) was used to obtain values for the qubit frequencies and anharmonicities. We defined a qubit as functional, when clear signatures are observed in one-tone and two-tone spectroscopy. With this definition, we determine the yield of our functional qubits to be 92 %.
@@ IV.2 Time domain characterization
To connect the wafer-level measurements at RT with the cryostatic qubit characterization, Fig. 8 plots the normal state resistance of a qubit’s Josephson Junction versus its frequency obtained by two-tone measurements at 10 mK. The plot includes the qubits presented in Fig. 7 (black squares) and data from two additional test wafers from different processing runs (green dots and blue triangles) to span a broader value range and check for reproducibility of the results. The data is fitted with Ambegaokar-Baratoff model [14]
with the Boltzmann constant , Planck constant , the electron charge , and the qubit capacitance being set constant to the design value of 86 fF, which was verified by evaluating the anharmonicity of the qubits. Leaving the critical temperature as a free fit parameter a value of T K is obtained, posing an offset of the curve from the ideal literature value of 1.2 K for aluminum. The inset of Fig. 8 shows a four-point DC measurement of an aluminum short on a test device with a channel width of 0.35 µm corresponding to the minimal dimension in our qubit device (width of bottom electrode). The measurement shows the onset of superconductivity at the expected 1.2 K, but full superconductivity is only reached at lower temperature of 0.82 K, which could explain the observed offset in the Ambegaokar-Baratoff relation shown in the main figure (Fig. 8). However, since this offset is stable between processing runs, the model can be used to predict the qubit frequency from the RT resistance measurement and hence poses a powerful tool towards precise frequency targeting by applying proper pre-selection of qubit chips.
@@ V Conclusion
In conclusion, this study presents a comprehensive methodology for the fabrication and characterization of superconducting qubits using a CMOS-compatible process on a 200 mm wafer. By integrating process control monitoring through the use of test structures, we successfully assessed the performance and variability of qubit junctions at RT, providing critical insights before cryogenic measurements. The analysis of shorts enabled an efficient evaluation of the metal structuring variability, leading to area variation of = 10.1 % and the calculated offset between design and fabrication width up to 99.3 nm. In combination with the junctions analysis of the resistance variation of = 12.4 %, we could calculate the oxide barrier variation to be = 7.2 %. Interestingly, we found that the breakthrough voltage has a similar across-wafer variation as the oxide barrier, making it highly important for detailed qubit RT characterization. In total, we achieved a yield of 92.8% for functional junctions across the 200 mm wafer. The cryogenic characterization demonstrated the effective targeting of qubit frequencies, with times reaching up to 80 µs and values exceeding 100 µs, indicative of high-quality devices. With the Ambegaokar-Baratoff formular applying to all our samples of several fabrication cycles with a 0.71 K, optimized frequency targeting was enabled. Our findings highlight the robustness and scalability of the fabrication process, positioning it as a viable pathway for future advancements in quantum processing units, particularly in the context of 3D integrated multi-chip systems [17, 18, 19]. Overall, this work lays the groundwork for the continued development and implementation of superconducting qubits in quantum computing applications.
@@ Acknowledgements
Moreover, we would like to thank Dr. Lars Nebrich for his support in layouting. We also appreciate Martin Heigl and Martin König for their help and valuable discussions in process development and the whole Fraunhofer EMFT clean room staff for the professional fabrication.

### v7

# CMOS-compatible processing and room-temperature characterization on wafer-level for scalable quantum computing

@@ Abstract
We report on an industry-grade CMOS-compatible qubit fabrication approach using a CMOS pilot line, enabling a yield of functional devices reaching 92.8 %, with a resistance spread evaluated across the full wafer 200 mm diameter of 12.4 % and relaxation times () approaching 80 µs. Furthermore, we conducted a comprehensive analysis of wafer-scale room temperature (RT) characteristics collected from multiple wafers and fabrication runs, focusing on RT measurements and their correlation to low temperature qubit parameters. From defined test structures, an across-wafer Josephson junction (JJ) area variation of 10.1 % and oxide barrier variation of 7.2 % was calculated. Additionally, from the room-temperature JJ characterization the qubit frequency can be derived on wafer-level applying the Ambegaokar-Baratoff model before low temperature measurements. This sets the stage for pre-cooldown wafer-level JJ evaluation and sorting. In particular, such early-on device characterization and validation are crucial for increasing the fabrication yield and qubit frequency targeting, which currently represent major scaling challenges. Furthermore, it enables the fabrication of large multichip quantum systems in the future. Our analysis highlight the great potential of CMOS-compatible industry-style fabrication of superconducting qubits for scalable quantum computing in a foundry pilot line cleanroom.
@@ I Introduction
Scalable, CMOS-compatible processing on 200 mm wafers is a critical step toward industrializing superconducting qubit-based quantum processing units [1]. It enables the precise fabrication and complex 3D integration needed for advanced quantum architectures [2, 3]. By leveraging mature semiconductor manufacturing methods, this approach ensures reproducibility, scalability, and the transition from laboratory prototypes to large-scale quantum systems. The fabrication of CMOS-compatible superconducting all-aluminum junctions using a double dry-etch technique with industry-standard manufacturing tools has been successfully demonstrated [4, 5, 6]. In addition, comprehensive wafer-scale characterization at RT has been reported [7, 8]. A recent study [4] integrated both approaches, showcasing the benefits of industry-compatible processes for transmon-style qubit fabrication, achieving a high yield of 98.5 % and relaxation times () of up to 167 µs. The study indicated that junction area and oxide barrier are essential parameters for the precise frequency targeting. As they vary across the wafer, they are worth monitoring for improved device characterisation. Additionally, like others [9, 10], they used critical dimension scanning electron microscopy (CD-SEM) to evaluate their structuring process at RT. Subsequently, they analyzed selected qubits after cooldown in the cryostat for benchmarking. This low temperature characterization, which is often very time-consuming and performed at die-level after leaving the production line, has become an established method for qubit characterization [11]. Currently, there is significant research focused on developing time-efficient, wafer-scale characterization methods for qubits during fabrication as part of the process control monitoring (PCM). This approach relies on RT techniques, as traditional cryogenic methods are impractical for large-scale analysis. Simple resistance measurements on test structures, like shorts and test junctions, can be used to monitor the line width variation as well as the junction oxide barrier variation across the 200 mm wafer. Additionally, a selection for functional qubit chips can be implemented to increase the yield for the subsequent process steps, as non-functional devices can be excluded early on with high probability. The primary aim of this wafer-scale characterization is to predict the cryogenic behavior of qubit devices, including the qubit frequency targeting, based on measurements conducted at RT.
@@
This present study demonstrates a comprehensive methodology for the fabrication and RT characterization of superconducting qubits on a 200 mm silicon wafer, emphasizing the importance of detailed wafer-level analysis for process monitoring. By integrating an array of test structures, including test junctions and shorts, our approach facilitates the precise evaluation of junction characteristics at RT and variation across the entire 200 mm wafer. For example, the analysis of shorts enabled an efficient evaluation of the metal structuring variability in terms of the junction area variation without using CD-SEM. Measuring the junctions resistance enabled an accurate selection of functional devices, which had a high yield across the full 200 mm wafer. In combination with the cryogenic analysis of those functional qubit chips, we could demonstrate validity of the Ambegaokar-Baratoff relation for all our samples over a large resistance range. This enables us to select qubit chips based on their resistance already at RT at wafer-level. Among others, this approach is especially important for further processing in 3D integrated architectures for increased yield. Our study highlights the effectiveness of the proposed characterization principles for CMOS-compatible superconducting qubits, which are crucial for achieving time-efficient process control and optimization for scaling of a QPU.
@@ II Methods
In total 375 identical dies with a size of 10 mm x 7 mm containing three different sub-dies were fabricated on a single wafer (Fig. 1a). The relevant sub-die, shown in Fig. 1b, is the qubit chip with a single reference resonator and four fixed-frequency qubits with equal Josephson junction (JJ) area, all addressed through a common RF feedline. Additionally, a process control monitoring (PCM) array was placed in the wafer’s dicing street. It consist of 16 test junctions of sizes varying from 0.1225 µ up to 1 µ and eight shorts (i.e. simple metal lines with landing pads for the prober needles) with widths from 0.35 µm to 1 µm. The fixed-frequency qubits were used to analyze the cryogenic performance of the fabrication process by measuring the qubits transition frequencies (), relaxation times (), and dephasing times (). With the test junctions, RT resistance and IV measurements were performed to access junction’s resistances and breakthrough voltage without harming the qubits JJ. Additionally, the shorts were used to obtain information about the JJ electrode width distribution across the 200 mm wafer, which is limited by the structuring process. Both test structures were used to maintain a time-efficient RT PCM during fabrication. Additionally, the sheet resistance of the bottom Al layer was measured using a standard four-point probe (4pp) setup to eliminate contact resistance. A standard 49-point sampling pattern across the 200 mm wafer was used to assess uniformity.
@@ III.1 Shorts Analysis
In general, for all our Al depositions for the layer thicknesses of 150 nm and 100 nm we measured of 1.6 %. As an example, Fig. 3c shows a 49-point standard 4pp measurement performed on a 200 mm wafer to determine the sheet resistance of a 150 nm thick Al film. Therefore, we assumed this value for the BE and TE deposition for the samples in this paper. For this, we obtain a value of = 5.1 %, which is smaller than the BE width distribution of = 8.7 %. Such increase in width variability for decreasing design width is frequently observed for metal structuring [10]. Additionally, a = 10.1 % for the area variation is obtained from the relation
The , observed for our junctions, with an area of 0.175 µm2, is notably larger than compared to values () reported sizes for JJs with areas (A 0.04 µm2) typically used in superconducting qubits [4, 10, 9]. When using optical lithography, the design widths approach the limits of the i-line stepper (365 nm) tool, which causes an increase in area variability. Using electron-beam lithography for junction patterning circumvents this limitation. We note that our time-efficient determination of electrode width variability, which applies only RT measurements of test structures, is much faster than the time-consuming optical inspection with critical dimension scanning electron microscope (CD-SEM) that is commonly used.
Furthermore, additional shorts with different widths can be used to determine the offset between the design width and actual width w. As the across-wafer resistance mean values of the respective shorts is inverse proportional to w, we can define . This approach can be used for both metal layers separately. The offset of the structuring process to the design value can be derived from Fig. 3b at the crossing of the fit with the y-axis and is = - 76.8 18.5 nm for the BE and = - 27.5 7.2 nm for the TE. These offsets result from approaching the feature-size limit of the optical lithography, which can be circumvented by electron-beam lithography as mentioned before. Thus, our PCM of shorts enable not only the determination of the area size variability across the wafer () in a time-efficient manner, but also provides a calculation of the offset from the design widths .
@@ III.2 Junction analysis
= 7.2 % is smaller than the area spread of = 10.1 %, indicating that the contribution of the spread in oxide thickness and barrier height is not dominant, although it remains larger than in related works [4]. By optimizing junction oxidation and Ar milling parameters, these variations can be further reduced. However, the miminum achievable resistance variation across the wafer needs to be further investigated. To further analyzing our junctions in terms of their tunnel barrier characteristics extended current voltage measurements were carried out. An example is shown in Fig. 5a. In general, the current through the junction can be described with a simple power-law dependence [15]
@@ IV.1 Frequency spectroscopy
The qubit characterization was conducted in accordance with well-established methodologies [17], [18]. The qubits under test were addressed using read-out resonators. By determining the “punch-out” power, i.e. the readout power at which the induced current exceeds the critical current of the junction, one-tone spectroscopy (Fig. 6a) serves as a first check whether an active qubit is coupled to the resonator. The strength of this coupling is indicated by magnitude of the shift. Two-tone spectroscopy (Fig. 6b) was used to obtain values for the qubit frequencies and anharmonicities. We defined a qubit as functional, when clear signatures are observed in one-tone and two-tone spectroscopy. With this definition, we determine the yield of our functional qubits to be 92 %.
@@ IV.2 Time domain characterization
To connect the wafer-level measurements at RT with the cryostatic qubit characterization, Fig. 8 plots the normal state resistance of a qubit’s Josephson Junction versus its frequency obtained by two-tone measurements at 10 mK. To confirm the reproducibility of the results, the plot includes the qubits presented in Fig. 7 (black squares) and additional data of two test wafers from independent fabrication runs (green dots and blue triangles), where changes to the Al etch chemistry lead to a shift in normal state resistance. This also spans the data set to a broader range for more accurate fitting applying the Ambegaokar-Baratoff model [14]
with the Boltzmann constant , Planck constant , the electron charge , and the qubit capacitance being set constant to the design value of 86 fF, which was verified by evaluating the anharmonicity of the qubits. Leaving the critical temperature as a free fit parameter, a value of T K is obtained, posing an offset of the curve from the expected literature value of 1.2 K for aluminum. The inset of Fig. 8 shows a four-point DC measurement of an aluminum short on a test device with a channel width of 0.35 µm corresponding to the minimal dimension in our qubit device (width of bottom electrode). The measurement shows the onset of superconductivity at the expected 1.2 K, but full superconductivity is only reached at a lower temperature of 0.82 K, which could explain the observed offset in the Ambegaokar-Baratoff relation shown in the main figure (Fig. 8). Furthermore, the shift could also origin from an effect offsetting the normal state resistance of the JJ like e.g. leakage currents through the substrate or other additional parasitic resistances within the current path. However, since this offset is stable between processing runs, the model can be used to predict the qubit frequency from the RT resistance measurement and hence poses a powerful tool towards precise frequency targeting by applying proper pre-selection of qubit chips.
@@ V Conclusion
In conclusion, this study presents a comprehensive methodology for the fabrication and characterization of superconducting qubits using a CMOS-compatible process on a 200 mm wafer. By integrating process control monitoring through the use of test structures, we successfully assessed the electrical characteristics and variability of qubit Josephson junctions at RT, providing critical insights before cryogenic measurements. The analysis of shorts enabled an efficient evaluation of the metal structuring variability, leading to area variation of = 10.1 % and the calculated offset between design and fabrication width up to 99.3 nm. In combination with the junctions analysis of the resistance variation of = 12.4 %, we could calculate the oxide barrier variation to be = 7.2 %. Interestingly, we found that the breakthrough voltage has a similar across-wafer variation as the oxide barrier, making it highly important for detailed qubit RT characterization. In total, we achieved a yield of 92.8% for functional junctions across the 200 mm wafer. The cryogenic characterization demonstrated the effective targeting of qubit frequencies, with times reaching up to 80 µs and values exceeding 100 µs, indicative of high-quality devices. With the Ambegaokar-Baratoff formular applying to all our samples of several fabrication cycles with a 0.71 K, optimized frequency targeting was enabled. Our results highlight the robustness and scalability of the fabrication process, positioning it as a viable pathway for future advancements in quantum processing units, particularly in the context of 3D integrated multi-chip systems [19, 20, 21]. Overall, this work lays the groundwork for the continued development and implementation of superconducting qubits in quantum computing applications.
@@ Acknowledgements
Moreover, we would like to thank Dr. Lars Nebrich for his support in layouting. We also appreciate Martin Heigl and Martin König for their help and valuable discussions in process development and the whole Fraunhofer EMFT clean room staff for the professional fabrication. We would also like to thank Samuel Taubenberger for supporting us in setting up the measurement software.